mads1309:

A tree 7 ft tall grows an average of 6 in. each year. Which equation models the tree’s height h after x years?

Extrinix:

Ok so, you know what the function of a linear equation is correct?

Extrinix:

\(y=mx+b\)

Extrinix:

So if a tree is already 7 feet tall, your b would be 7. \(y=mx+7\) Then, it says it grows by 6 inches a year, so when converted to feet, you would have \(\dfrac{1}{2}\). \(y=\dfrac{1}{2}x+7\) So that would be your equation. \(y=\dfrac{1}{2}x+7\)